### v1

# 4.2cm4.2cm3.6cm3cm

Abstract: In this paper, generalized Loop Heisenberg-Virasoro algebras are introduced. Firstly, we determine the derivations on generalized Loop Heisenberg-Virasoro algebras. Then we show that all $2$-local derivations are derivations. Furthermore, we determine the biderivations on generalized Loop Heisenberg-Virasoro algebras are inner biderivations and give their applications. Finally, the automorphism groups on generalized Loop Heisenberg-Virasoro algebras are presented.
Keywords: generalized Loop Heisenberg-Virasoro algebras; derivations; $2$-local derivations; biderivations; automorphism groups
Mathematics Subject Classification 2020: 17B40, 17B65, 17B68
Introduction
The Heisenberg-Virasoro algebra is an important infinite dimensional Lie algebra and it plays a vital role in numerous fields of mathematics and physics. The Heisenberg-Virasoro algebra is the universal central extension of the Lie algebra of differential operators of degree at most 1 defined on the circle, and has been extensively studied in mathematical and physical literature.
As is well known, structure theory is essential in the research of Lie algebras. Likewise, the derivation algebra and automorphism group on the generalized Loop Virasoro algebra, the generalized Loop Schrödinger–Virasoro algebras and the generalized Heisenberg-Virasoro algebra, etc. have been determined respectively (see [6, 7, 9, 10, 19]). $2$-local derivations on the generalized Witt algebras are also obtained in [5]. The author of [20] proved that all 2-local derivations on the Witt algebra as well as on the positive Witt algebra are derivations, and give an example of infinite-dimensional Lie algebra with a 2-local derivation which is not a derivation. The author of [3] determined biderivations on $m_{0},\,L_{1},\,m_{2}$ which are three Lie algebras of maximal class with one-dimensional homogeneous components, and characterized their local and 2-local derivations. All biderivations on the twisted Heisenberg–Virasoro algebra, the Schrödinger-Virasoro algebra and the W-algebras, etc. have been given in [11, 12, 13, 18].
Likewise, representation theory is also very important in the research of Lie algebras. For example, the author of [2] described the structure of the irreducible highest weight modules for the twisted Heisenberg-Virasoro Lie algebra at level zero, sufficient and necessary conditions are given for these non-weight modules on the mirror Heisenberg-Virasoro algebra to be irreducibie (see [14]), the author of [15, 16, 17] also investigated representations on the mirror Heisenberg-Virasoro algebra.
Inspired by [8], we introduced the generalized Loop Heisenberg-Virasoro algebra. When $\Gamma=\mathbb{Z}$, the generalized Loop Heisenberg-Virasoro algebra in the paper is the Loop Heisenberg-Virasoro algebra [8]. The aim is to determine the derivations, $2$-local derivations, biderivations and automorphism groups of generalized Loop Heisenberg-Virasoro algebra by use of the methods for derivations, $2$-local derivations, biderivations, and automorphism groups of respective Lie algebras in [7, 9, 12, 20]. The structure of this paper is as follows: Firstly, based on its relevant properties, the derivations on generalized Loop Heisenberg-Virasoro algebra are determined. On this basis, it is proved that all $2$-local derivations of this algebra are derivations. Secondly, based on the fact that it is a perfect Lie algebra, we conclude that all biderivations are inner biderivations. Then we prove the necessary and sufficient conditions for its commutating linear maps and any commutative $post$-Lie algebra structure on this algebra is trivial. Finally, we define five types of automorphism groups on it and prove that any automorphism group of the algebra can be represented the direct product of the five types of automorphisms.
1. Preliminaries
Throughout this paper, the ground field $\mathbb{F}$ is an algebraically closed field of characteristic zero and all vector spaces, algebras are over $\mathbb{F}$. $\Gamma$ represents an abelian additive group, and $\mathbb{Z}$ represents the set of integers. Let $\mathrm{span}\{X\}$ denote the vector space generated by the vectors in $\{X\}$.
Definition 1.
Generalized Loop Heisenberg-Virasoro algebra is a Lie algebra with basis $\left\{L_{\alpha,i},H_{\beta,j}\,|\,\alpha,\beta\in\Gamma,\,i,j\in\mathbb{Z}\right\}$, subject to the following nontrivial Lie brackets:
$[L_{\alpha,i},L_{\beta,j}]$ $=$ $(\alpha-\beta)L_{\alpha+\beta,i+j},$
$[L_{\alpha,i},H_{\beta,j}]$ $=$ $-\beta H_{\alpha+\beta,i+j},$
$[H_{\alpha,i},H_{\beta,j}]$ $=$ $0$
In the sequel, generalized Loop Heisenberg-Virasoro algebra denoted by $\mathcal{L}(\Gamma)$. It is worth noting that Loop Heisenberg-Virasoro algebra in [8] is exactly $\mathcal{L}(\Gamma)$ when $\mathbb{Z}=\Gamma$. Let $\mathcal{LV}$ be subalgebra of $\mathcal{L}(\Gamma)$ spanned by $\{L_{\alpha,i}\,|\,\alpha\in\Gamma,\,i\in\mathbb{Z}\}$. Then $\mathcal{LV}$ is actually isomorphic to the centerless generalized Loop Virasoro algebra (see [7]). In addition, $\mathcal{L}(\Gamma)$ contains the generalized Heisenberg-Virasoro algebra
$\mathcal{HV}=\mathrm{span}_{\mathbb{F}}\{L_{\alpha,0},H_{\beta,0}\,|\,\alpha,\beta\in\Gamma\}.$
Let $L$ be a Lie algebra and $V$ be $L$-module. A linear map $\mathcal{D}:L\rightarrow V$ is called a derivation of $\mathcal{L}(\Gamma)$ if
$\mathcal{D}([x,y])=x\cdot\mathcal{D}(y)+y\cdot\mathcal{D}(x),\mbox{for all}\,x,y\in L.$
A linear map $\mathrm{ad}:x\mapsto x\cdot v$ is called the inner derivation. Otherwise others are called outer derivations.
Let $\mathrm{Der}(\mathcal{L}(\Gamma))$ and $\mathrm{Inn}(\mathcal{L}(\Gamma))$ be the vector spaces consisting of all derivations and all inner derivations on $\mathcal{L}(\Gamma)$, respectively. Then
$H^{1}(\mathcal{L}(\Gamma))\cong\mathrm{Der}(\mathcal{L}(\Gamma))/\mathrm{Inn}(\mathcal{L}(\Gamma))$
is the first cohomology group of $\mathcal{L}(\Gamma)$ (see [9]).
Since $L_{0,0}$ is semisimple, then the $\Gamma$-grading on $\mathcal{L}(\Gamma)$ is given,
$\mathcal{L}(\Gamma)=\bigoplus_{\mu\in\Gamma}\mathcal{L}(\Gamma)_{\mu},$
where
$\mathcal{L}(\Gamma)_{\mu}=\left\{x\in\mathcal{L}(\Gamma)\,|\,[L_{0,0},x]=-\mu x\right\}=\left\{L_{\mu,j},H_{\mu,j}\,|\,j\in\mathbb{Z}\right\}.$
Recall that a derivation $\mathcal{D}\in\mathrm{Der}\mathcal{L}(\Gamma)$ is of degree $\gamma\in\Gamma$ if $D(\mathcal{L}(\Gamma)_{\alpha})\subset\mathcal{L}_{\alpha+\gamma}$ for all $\alpha\in\Gamma$. Let $(\mathrm{Der}\mathcal{L}(\Gamma))_{\gamma}$ be the space of all derivations of degree $\gamma$.
2. Derivations on $\mathcal{L}(\Gamma)$
Lemma 2.
$\mathrm{Der}\mathcal{L}(\Gamma)=(\mathrm{Der}\mathcal{L}(\Gamma))_{0}\bigoplus\mathrm{ad}\mathcal{L}(\Gamma)$.
Proof.
(i) For all $\mathcal{D}\in\mathrm{Der}\mathcal{L}(\Gamma),x_{\alpha}\in\mathcal{L}(\Gamma)_{\alpha}$, assume that $\mathcal{D}(x_{\alpha})=\sum_{\beta\in\Gamma}y_{\beta}$. We define $\mathcal{D}_{\gamma}(x_{\alpha})=y_{\alpha+\gamma}$, where $\mathcal{D}_{\gamma}\in(\mathrm{Der}\mathcal{L}(\Gamma))_{\gamma}$. It is easy to prove that
$\mathcal{D}_{\gamma}([x_{\alpha},x_{\beta}])=[\mathcal{D}_{\gamma}(x_{\alpha}),x_{\beta}]+[x_{\alpha},\mathcal{D}_{\gamma}(x_{\beta})],\,\forall x_{\alpha}\in\mathcal{L}(\Gamma)_{\alpha},x_{\beta}\in\mathcal{L}(\Gamma)_{\beta}.$
Let $x=\sum_{\alpha\in\Gamma}x_{\alpha}$, where $x_{\alpha}\in\mathcal{L}(\Gamma)_{\alpha}$. We can conclude that
$\mathcal{D}(x)=\sum_{\alpha\in\Gamma}\sum_{\gamma\in\Gamma}\mathcal{D}_{\gamma}(x_{\alpha})=\sum_{\gamma\in\Gamma}\mathcal{D}_{\gamma}(x).$
That is for any derivation $\mathcal{D}\in\mathcal{L}(\Gamma)$, one has
$\mathcal{D}=\sum_{\gamma\in\Gamma}\mathcal{D}_{\gamma},\,\mathcal{D}_{\gamma}\in(\mathrm{Der}\mathcal{L}(\Gamma))_{\gamma}.$
such that for all $x\in\mathcal{L}(\Gamma)$, only finitely $\mathcal{D}_{\gamma}(x)\neq 0$, and $\mathcal{D}(x)=\sum_{\gamma\in\Gamma}\mathcal{D}_{\gamma}(x)$.
(ii) For all $x\in\mathcal{L}(\Gamma)$, one hand it has
$\mathcal{D}([L_{0,0},x])$ $=$ $[\mathcal{D}(L_{0,0}),x]+[L_{0,0},\mathcal{D}(x)]$
$=$ $[\mathcal{D}(L_{0,0}),x]+[L_{0,0},y_{\mu+\gamma}]$
$=$ $[\mathcal{D}(L_{0,0}),x]+{-(\mu+\gamma)}\mathcal{D}(x)$,
other hand
$\mathcal{D}([L_{0,0},x])$ $=$ $\mathcal{D}(-\mu x)$
$=$ $-\mu\mathcal{D}(x)$,
then
$\mathcal{D}(x)$ $=$ $[\frac{1}{\gamma}\mathcal{D}(L_{0,0}),x]$
$=$ ${\rm{ad}}_{\gamma^{-1}\mathcal{D}(L_{0,0})}(x)$.
Thus $\mathcal{D}=\mathrm{ad}_{\gamma^{-1}\mathcal{D}(L_{0,0})}$ is an inner derivation. When $\gamma\neq 0$, for all $\mathcal{D}\in(\mathrm{Der}\mathcal{L}(\Gamma))_{\gamma}$ is an inner derivation.
(iii) For all $\mathcal{D}\in(\mathrm{Der}\mathcal{L}(\Gamma))_{0}\cap\mathrm{ad}\mathcal{L}(\Gamma)$. On the one hand, there exists an $X\in\mathcal{L}(\Gamma)$ such that $\mathcal{D}(Y_{\alpha,i})=[X,Y_{\alpha,i}]$, where $Y\in\mathcal{L}(\Gamma),\alpha\in\Gamma,\,i\in\mathbb{Z}$.
On the other hand, by $\mathcal{D}\in(\mathrm{Der}\mathcal{L}(\Gamma))_{0}$, then for $Z_{\gamma,k}\in\mathcal{L}(\Gamma)_{\gamma}$, where $\gamma\in\Gamma,k\in\mathbb{Z}$, we have that $\mathcal{D}(Z_{\gamma,k})=[X,Z_{\gamma,k}]\in\mathcal{L}(\Gamma)_{\gamma}$.
By $X=\sum_{\begin{subarray}{c}\beta\in\Gamma\\ j\in\mathbb{Z}\end{subarray}}X_{\beta,j}$, where $X_{\beta,j}\in\mathcal{L}(\Gamma)_{\beta}$, we have
$[X,Z_{\gamma,k}]=\sum_{\begin{subarray}{c}\beta\in\Gamma\\ j\in\mathbb{Z}\end{subarray}}[X_{\beta,j},Z_{\gamma,k}]\in\mathcal{L}(\Gamma)_{\gamma}.$
Thus $[X_{\beta,j},Z_{\gamma,k}]=0,\,\forall\beta\in\Gamma,\,j\in\mathbb{Z}$, that is $\mathcal{D}=0$. ∎
By Lemma 2, if we want to determine $\mathrm{Der}\mathcal{L}(\Gamma)$, only needs to define $(\mathrm{Der}\mathcal{L}(\Gamma))_{0}$. For all $\phi\in\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t^{1},t^{-1}]),g\in\mathbf{g}(\Gamma),b\in\mathbb{F}[t^{1},t^{-1}],\rho\in\mathbb{F}[t^{1},t^{-1}]\frac{d}{dt}$ and
$\mathbf{g}(\Gamma)=\left\{g:\Gamma\rightarrow\mathbb{F}[t^{1},t^{-1}]\,|\,(\alpha-\beta)g_{\alpha+\beta}=\alpha g_{\alpha}-\beta g_{\beta},\,\forall\alpha,\beta\in\Gamma\right\},$
where $\mathbb{F}[t^{1},t^{-1}]\frac{d}{dt}$ is a derivation algebra of $\mathbb{F}[t^{1},t^{-1}]$. We brief note $\mathbf{g}(\Gamma)$ by $\mathbf{g}$ in the sequel. The following linear maps $D_{\phi},\,D_{g},\,D_{b},\,D^{\rho}$ on $\mathcal{L}(\Gamma)$ are given by respectively:
$D_{\phi}(X_{\mu,i})$ $=$ $\phi(\mu)X_{\mu,i}$,
$D_{g}(L_{\alpha,i})$ $=$ $g_{\alpha}H_{\alpha,i},\quad D_{g}(H_{\alpha,i})=0$,
$D_{b}(L_{\alpha,i})$ $=$ $0,\quad D_{b}(H_{\alpha,i})=bH_{\alpha,i}$,
$D^{\rho}(X_{\mu,i})$ $=$ $X_{\mu}\rho(t^{i}),\,\forall X\in\{L,H\},\,i\in\mathbb{Z},\,\mu,\alpha\in\Gamma$.
It is easy to see that these operators defined above are all homogeneous derivations of degree zero. Let
$\mathcal{D}_{\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t^{1},t^{-1}])}\,:=\,\left\{D_{\phi}\,|\,\phi\in\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t^{1},t^{-1}])\right\}$,
$\mathcal{D}_{\mathbf{g}(\Gamma)}\,:=\,\left\{D_{g}\,|\,g\in\mathbf{g}(\Gamma)\right\}$,
$\mathcal{D}_{\mathbb{F}[t^{1},t^{-1}]}\,:=\,\left\{D_{b}\,|\,b\in\mathbb{F}[t^{1},t^{-1}]\right\}$,
$\mathcal{D}_{\mathbb{F}[t^{1},t^{-1}]\frac{d}{dt}}\,:=\,\{D^{\rho}\,|\,\rho\in\mathbb{F}[t^{1},t^{-1}]\frac{d}{dt}\}$.
Lemma 3.
Let $D\in(\mathrm{Der}\mathcal{L}(\Gamma))_{0}$. Then there exists $f\in\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t^{1},t^{-1}]),\,g\in\mathbf{g}$ such that
$D(L_{\alpha,i})=f_{\alpha}L_{\alpha,i}+g_{\alpha}H_{\alpha,i},\,\alpha\in\Gamma,i\in\mathbb{Z}.$
Proof.
Let $D(L_{\alpha,i})=f_{\alpha,i}L_{\alpha,i}+g_{\alpha,i}H_{\alpha,i},\,f_{\alpha,i},g_{\alpha,i}\in\mathbb{F}[t^{1},t^{-1}]$. We have
$D([L_{\alpha,i},L_{\beta,j}])$ $=$ $D({(\alpha-\beta)L_{\alpha+\beta,i+j}})$
$=$ $(\alpha-\beta)D(L_{\alpha+\beta,i+j})$
$=$ $(\alpha-\beta)(f_{\alpha+\beta,i+j}L_{\alpha+\beta,i+j}+g_{\alpha+\beta,i+j}H_{\alpha+\beta,i+j})$.
On the other hand,
$D([L_{\alpha,i},L_{\beta,j}])$ $=$ $[D(L_{\alpha,i}),L_{\beta,j}]+[L_{\alpha,i},D(L_{\beta,j})]$
$=$ $[f_{\alpha,i}L_{\alpha,i}+g_{\alpha,i}H_{\alpha,i},L_{\beta,j}]+[L_{\alpha,i},f_{\beta,j}L_{\beta,j}+g_{\beta,j}H_{\beta,j}]$
$=$ $(\alpha-\beta)f_{\alpha,i}L_{\alpha+\beta,i+j}+\alpha g_{\alpha,i}H_{\alpha+\beta,i+j}+(\alpha-\beta)f_{\beta,j}L_{\alpha+\beta,i+j}$
$-\beta g_{\beta,j}H_{\alpha+\beta,i+j}$.
We obtain
$f_{\alpha+\beta,i+j}=f_{\alpha,i}+f_{\beta,j},\,\alpha\neq\beta$
$(\alpha-\beta)g_{\alpha+\beta,i+j}=\alpha g_{\alpha,i}-\beta g_{\beta,j},\,\forall\alpha,\beta\in\Gamma,\,i,j\in\mathbb{Z}$.
Let $\alpha=0,i=0$. Then
$f_{0+\beta,0+j}=f_{0,0}+f_{\beta,j}.$
Thus $f_{0,0}=0$, Let $\beta=-\alpha\neq 0,i=j=0$. Then
$0=f_{0,0}=f_{\alpha,0}+f_{-\alpha,0},\,\forall\alpha\neq 0$
Let $0\neq\eta\in\Gamma$. Satisfying $\eta\neq\alpha,\,-\eta\neq\beta,\alpha+\eta\neq\beta-\alpha$, then
$f_{\alpha+\beta,i+j}=f_{\alpha+\eta,i}+f_{\beta-\eta,j}$,
$f_{\alpha+\eta,i}=f_{\alpha,i}+f_{\eta,0}$,
$f_{\beta-\eta.j}=f_{\beta,j}+f_{-\eta,0}$.
That is
$f_{\alpha+\beta,i+j}=f_{\alpha,i}+f_{\beta,j},\,\forall\alpha,\beta\in\Gamma,i,j\in\mathbb{Z}.$
Then $f_{\alpha,i}=f_{\alpha,0}+f_{0,i}$. Let $\rho=tf_{0,1}\frac{d}{dt}$. We have
$D^{\rho}(L_{0,1})=L_{0}(tf_{0,1}\frac{d}{dt})(t)=L_{0}tf_{0,1}.$
Replace $D$ by $D-D^{\rho}$. Let $D(L_{0,1})=0$. Then $f_{0,1}=0$ and $f_{0,i}=if_{0,1}=0$, thus
$f_{\alpha,i}=f_{\alpha,0}+f_{0,i}=f_{\alpha,0},\,\forall\alpha\in\Gamma,i\in\mathbb{Z}.$
For convenience, denote $f_{\alpha,i}$ as $f_{\alpha}$, then
$f_{\alpha+\beta}=f_{\alpha}+f_{\beta},\,f\in\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t^{1},t^{-1}]).$
Let $\alpha=0,j=0$. We have
$g_{\beta,i}=g_{\beta,0}.$
Let $\alpha=-\beta\neq 0$. We have
$2\beta g_{0,i+j}$ $=$ $\beta g_{\beta,j}+\beta g_{-\beta,i}$
$=$ $\beta g_{\beta,0}+\beta g_{-\beta,0}$
$=$ $2\beta g_{0,0}$
$=$ 00.
Thus
$g_{\beta,i}=g_{\beta,0}-g_{0,i}=g_{\beta,0},\,\forall\beta\in\Gamma,i\in\mathbb{Z}.$
For convenience, denote $g_{\beta}$ as $g_{\beta,i}$, then
$(\alpha-\beta)g_{\alpha+\beta}=\alpha g_{\alpha}-\beta g_{\beta},\,\forall g\in\mathbf{g}(\Gamma).$
Lemma 4.
Let $D\in(\mathrm{Der}\mathcal{L}(\Gamma))_{0}$. Then there exists the $b\in\mathbb{F}[t^{1},t^{-1}]$ such that
$D(H_{\alpha,i})=(f_{\alpha}+b)H_{\alpha,i},\,\alpha\in\Gamma,i\in\mathbb{Z}.$
Proof.
For all $\alpha\in\Gamma,i\in\mathbb{Z}$, let $D(H_{\alpha,i})=\phi_{\alpha,i}L_{\alpha,i}+\psi_{\alpha,i}H_{\alpha,i},\,\phi_{\alpha,i},\psi_{\alpha,i}\in\mathbb{F}[t^{1},t^{-1}].$
$D([L_{-\alpha,-i},H_{\alpha,i}])$ $=$ $D(-\alpha H_{0,0})$
$=$ $-\alpha(\phi_{0,0}L_{0,0}+\psi_{0,0}H_{0,0})$.
On the other hand,
$D([L_{-\alpha,-i},H_{\alpha,i}])$ $=$ $[D(L_{-\alpha,-i}),H_{\alpha,i}]+[L_{-\alpha,-i},D(H_{\alpha,i})]$
$=$ $[f_{-\alpha}L_{-\alpha,-i}+g_{-\alpha}H_{-\alpha,-i},H_{\alpha,i}]+[L_{-\alpha,-i},\phi_{\alpha,i}L_{\alpha,i}+\psi_{\alpha,i}H_{\alpha,i}]$
$=$ $-\alpha f_{-\alpha}H_{0,0}-2\alpha\phi_{\alpha,i}L_{0,0}-\alpha\phi_{\alpha,i}H_{0,0}$.
We have $\alpha(2\psi_{\alpha,i}-\psi_{0,0})L_{0,0}+\alpha(f_{-\alpha}+\phi_{\alpha,i}-\phi_{0,0})H_{0,0}=0$, thus
$\phi_{\alpha,i}=\frac{1}{2}\phi_{0,0},\,f_{-\alpha}+\psi_{\alpha,i}=\psi_{0,0},\,\forall 0\neq\alpha\in\Gamma.$
Since $H_{0,0}\in C(\mathcal{L}(\Gamma))$, then $D(H_{0,0})\in C(\mathcal{L}(\Gamma))$, thus $\phi_{0,0}=0$, so $\phi_{\alpha,i}=0,\,\forall\alpha\in\Gamma,i\in\mathbb{Z}$, we have
$D(H_{\alpha,i})$ $=$ $\psi_{\alpha,i}H_{\alpha,i}$
$=$ $(\psi_{0,0}-f_{-\alpha})H_{\alpha,i}$
$=$ $(f_{\alpha}+\psi_{0,0})H_{\alpha,i}$.
Let $\psi_{0,0}=b$. Then $D(H_{\alpha,i})=(f_{\alpha}+b)H_{\alpha,i}$. ∎
Theorem 5.
$\mathrm{Der}\mathcal{L}(\Gamma)=\mathrm{ad}(\mathcal{L}(\Gamma))\bigoplus(\mathcal{D}_{\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t^{1},t^{-1}])}\oplus\mathcal{D}_{\mathbf{g}(\Gamma)}\oplus\mathcal{D}_{\mathbb{F}[t^{1},t^{-1}]}\oplus\mathcal{D}_{\mathbb{F}[t^{1},t^{-1}]\frac{d}{dt}})$.
Proof.
(i)$(\mathrm{Der}\mathcal{L}(\Gamma))_{0}=\mathcal{D}_{\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t^{1},t^{-1}])}+\mathcal{D}_{\mathbf{g}(\Gamma)}+\mathcal{D}_{\mathbb{F}[t^{1},t^{-1}]}+\mathcal{D}_{\mathbb{F}[t^{1},t^{-1}]\frac{d}{dt}}$.
In fact, by Lemma 3 and Lemma 4, for all $D\in\mathcal{D}_{\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t^{1},t^{-1}])}$, there exists $\rho\in\mathbb{F}[t^{1},t^{-1}]\frac{d}{dt},f\in\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t^{1},t^{-1}]),g\in\mathbf{g}(\Gamma),b\in\mathbb{F}[t^{1},t^{-1}]$, such that
$(D-D^{\rho})(L_{\alpha,i})$ $=$ $f_{\alpha}L_{\alpha,i}+g_{\alpha}H_{\alpha,i}$,
$(D-D^{\rho})(H_{\alpha,i})$ $=$ $(f_{\alpha}+b)H_{\alpha,i}$.
On the other hand, it follows from the definition of $D_{f},D_{g}$ and $D_{b}$ that
$(D_{f}+D_{g}+D_{b})(L_{\alpha,i})$ $=$ $f_{\alpha}L_{\alpha,i}+g_{\alpha}H_{\alpha,i}$,
$(D_{f}+D_{g}+D_{b})(H_{\alpha,i})$ $=$ $(f_{\alpha}+b)H_{\alpha,i}$.
Thus $D=D^{\rho}+D_{f}+D_{g}+D_{b}$.
(ii)$\mathcal{D}_{\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t^{1},t^{-1}])}+\mathcal{D}_{\mathbf{g}(\Gamma)}+\mathcal{D}_{\mathbb{F}[t^{1},t^{-1}]}+\mathcal{D}_{\mathbb{F}[t^{1},t^{-1}]\frac{d}{dt}}$ is a direct sum.
In fact, let
$D=D_{\phi}+\sum_{j\in J}c_{j}D_{g_{j}}+D_{b}+D^{\rho}.$
$J$ is a finite index set, $\phi\in\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t^{1},t^{-1}]),\,g_{j}\in\mathbf{g}(\Gamma),\,b\in\mathbb{F}[t^{1},t^{-1}],\,\rho\in\mathbb{F}[t^{1},t^{-1}]\frac{d}{dt},\,c_{j}\in\mathbb{F}$. If $D=0$, to prove that $D_{\phi}=\sum_{j\in J}c_{j}D_{g_{j}}=D_{b}=D^{\rho}=0$, we have
$D(L_{\alpha,i})$ $=$ $0=f_{\alpha}L_{\alpha,i}+g_{\alpha}H_{\alpha,i},$
$D(H_{\alpha,i})$ $=$ $0=(f_{\alpha}+b)H_{\alpha,i}.$
On the other hand,
$(D_{\phi}+\sum_{j\in J}c_{j}D_{g_{j}}+D_{b}+D^{\rho})(L_{\alpha,i})$ $=$ 00,
$(D_{\phi}+\sum_{j\in J}c_{j}D_{g_{j}}+D_{b}+D^{\rho})(H_{\alpha,i})$ $=$ 00.
Then $D_{\phi}=\sum_{j\in J}c_{j}D_{g_{j}}=D_{b}=D^{\rho}=0$. ∎
3. $2$-local derivations on $\mathcal{L}(\Gamma)$
Recall that a map $\Delta:\mathcal{L}\rightarrow\mathcal{L}$ (not linear in general) is called a $2$-local derivation if for every $x,y\in\mathcal{L}$, there exists a derivation $\Delta_{x,y}:\mathcal{L}\rightarrow\mathcal{L}$ such that $\Delta(x)=\Delta_{x,y}(x)$ and $\Delta(y)=\Delta_{x,y}(y)$ (see [4]).
By Theorem 5, we have the following result.
Corollary 6.
Let $\Delta$ be a $2$-local derivation on $\mathcal{L}(\Gamma)$. For all $x,y\in\mathcal{L}(\Gamma)$, there exist derivations $\Delta_{x,y}$ on $\mathcal{L}(\Gamma)$, such that $\Delta(x)=\Delta_{x,y}(x)\,,\,\Delta(y)=\Delta_{x,y}(y)$, and it can be written as
$\Delta_{x,y}=\mathrm{ad}(\sum_{\begin{subarray}{c}\alpha\in\Gamma\\ i\in\mathbb{Z}\end{subarray}}(a_{\alpha,i}(x,y)L_{\alpha,i}+b_{\alpha,i}(x,y)H_{\alpha,i}))+\lambda_{1}(x,y)D_{f}+\lambda_{2}(x,y)D_{g}+\lambda_{3}(x,y)D_{b}+\lambda_{4}(x,y)D^{\rho}.$
where $a_{\alpha,i},b_{\alpha,i},\lambda_{i}(i=1,2,3,4)$ are complex-valued functions on $\mathcal{L}(\Gamma)\times\mathcal{L}(\Gamma)$, $D_{f},D_{g},D_{b},D^{\rho}$ is given previous.
Lemma 7.
Let $\Delta$ be a $2$-local derivation on $\mathcal{L}(\Gamma)$. For any fixed $x\in\mathcal{L}(\Gamma)$, if $\Delta(L_{\alpha,i})=0$, $\forall\alpha\in\Gamma,i\in\mathbb{Z}$, then
$\Delta_{L_{\alpha,i},x}$ $=$ $\mathrm{ad}(\sum_{\begin{subarray}{c}\alpha\in\Gamma\\ j\in\mathbb{Z}\end{subarray}}(a_{\alpha,j}(L_{\alpha,i},x)L_{\alpha,j}+b_{0,j}(L_{\alpha,i},x)H_{0,j}))+(1-\delta_{\alpha,0})\lambda_{1}(L_{\alpha,i},x)D_{f}$
$+(1-\delta_{\alpha,0})\lambda_{2}(L_{\alpha,i},x)D_{g}+\lambda_{3}(L_{\alpha,i},x)D_{b}+(1-\delta_{i,0})\lambda_{4}(L_{\alpha,i},x)D^{\rho}$.
Proof.
If $\Delta(L_{\alpha,i})=0$, we have
$\Delta(L_{\alpha,i})$ $=$ $\Delta_{L_{\alpha,i},x}(L_{\alpha,i})$
$=$ $[\sum_{\begin{subarray}{c}\alpha\in\Gamma\\ j\in\mathbb{Z}\end{subarray}}(a_{\beta,j}(L_{\alpha,i},x)L_{\beta,j}+b_{\beta,j}(L_{\alpha,i},x)H_{\beta,j}),L_{\alpha,i}]+\lambda_{1}(L_{\alpha,i},x)D_{f}(L_{\alpha,i})$
$+\lambda_{2}(L_{\alpha,i},x)D_{g}(L_{\alpha,i})+\lambda_{3}(L_{\alpha,i},x)D_{b}(L_{\alpha,i})+\lambda_{4}(L_{\alpha,i},x)D^{\rho}(L_{\alpha,i})$
$=$ $\sum_{\begin{subarray}{c}\alpha\in\Gamma\\ j\in\mathbb{Z}\end{subarray}}((\beta-\alpha)a_{\beta,j}(L_{\alpha,i},x)L_{\alpha+\beta,i+j}+\beta b_{\beta,j}(L_{\alpha,i},x)H_{\alpha+\beta,i+j})$
$+f_{\alpha}\lambda_{1}(L_{\alpha,i},x)L_{\alpha,i}+g_{\alpha}\lambda_{2}(L_{\alpha,i},x)L_{\alpha,i}+i\lambda_{4}(L_{\alpha,i},x)L_{\alpha,i-1}$
$=$ 00.
Thus $a_{\beta,j}(L_{\alpha,i},x)=0\,(\alpha\neq\beta),\,b_{\beta,j}(L_{\alpha,i},x)=0\,(\beta\neq 0),\,\lambda_{1}(L_{0,i},x)=\lambda_{2}(L_{0,i},x)=\lambda_{4}(L_{\alpha,0},x)=0.$
Lemma 8.
Let $\Delta$ be a $2$-local derivation on $\mathcal{L}(\Gamma)$ such that $\Delta(L_{0,i})=\Delta(L_{1,j})=0$. Then $\Delta(L_{\alpha,k})=0,\,\forall\alpha\in\Gamma,\,k\in\mathbb{Z}$.
Proof.
Since $\Delta(L_{0,i})=\Delta(L_{1,j})=0$, we can assume
$\Delta_{L_{\gamma,k},x}$ $=$ $\mathrm{ad}(a_{\gamma,j}(L_{\gamma,k},x)L_{\gamma,j}+b_{0,j}(L_{\gamma,k},x)H_{0,j})+(1-\delta_{\gamma,0})\lambda_{1}(L_{\gamma,k},x)D_{f}$
$+(1-\delta_{\gamma,0})\lambda_{2}(L_{\gamma,k},x)D_{g}+\lambda_{3}(L_{\gamma,k},x)D_{b}+(1-\delta_{k,0})\lambda_{4}(L_{\gamma,k},x)D^{\rho},\,\gamma=0,1,\,\forall x\in\mathcal{L}(\Gamma)$.
Let $x=L_{\alpha,k}$. We have
$\Delta(L_{\alpha,k})$ $=$ $\Delta_{L_{0,i},L_{\alpha,k}}(L_{\alpha,k})$
$=$ $[a_{0,z}(L_{0,i},L_{\alpha,k})L_{0,z}+b_{0,z}(L_{0,i},L_{\alpha,k})H_{0,z},L_{\alpha,k}]+(1-\delta_{0,0})\lambda_{1}(L_{0,i},L_{\alpha,k})D_{f}(L_{\alpha,k})$
$+(1-\delta_{0,0})\lambda_{2}(L_{0,i},L_{\alpha,k})D_{g}(L_{\alpha,k})+\lambda_{3}(L_{0,i},L_{\alpha,k})D_{b}(L_{\alpha,k})$
$+(1-\delta_{i,0})\lambda_{4}(L_{0,i},L_{\alpha,k})D^{\rho}(L_{\alpha,k})$
$=$ $-\alpha a_{0,z}(L_{0,i},L_{\alpha,k})L_{\alpha,z+k}+k(1-\delta_{i,0})\lambda_{4}(L_{0,i},L_{\alpha,k})L_{\alpha,k-1}$,
$\Delta(L_{\alpha,k})$ $=$ $\Delta_{L_{1,j},L_{\alpha,k}}(L_{\alpha,k})$
$=$ $[a_{0,z}(L_{1,j},L_{\alpha,k})L_{1,z}+b_{0,z}(L_{1,j},L_{\alpha,k})H_{0,z},L_{\alpha,k}]+(1-\delta_{1,0})\lambda_{1}(L_{1,j},L_{\alpha,k})D_{f}(L_{\alpha,k})$
$+(1-\delta_{1,0})\lambda_{2}(L_{1,j},L_{\alpha,k})D_{g}(L_{\alpha,k})+\lambda_{3}(L_{1,j},L_{\alpha,k})D_{b}(L_{\alpha,k})$
$+(1-\delta_{j,0})\lambda_{4}(L_{1,j},L_{\alpha,k})D^{\rho}(L_{\alpha,k})$
$=$ $(1-\alpha)a_{1,z}(L_{1,j},L_{\alpha,k})L_{\alpha+1,z+k}+f_{\alpha}\lambda_{1}(L_{1,j},L_{\alpha,k})L_{\alpha,k}+g_{\alpha}\lambda_{2}(L_{1,j},L_{\alpha,k})L_{\alpha,k}$
$+k(1-\delta_{j,0})\lambda_{4}(L_{1,j},L_{\alpha,k})L_{\alpha,k-1}$.
Comparing the formulas above, we have
$a_{0,z}(L_{0,i},L_{\alpha,k})=a_{1,z}(L_{1,j},L_{\alpha,k})=\lambda_{1}(L_{1,j},L_{\alpha,k})=\lambda_{2}(L_{1,j},L_{\alpha,k})=0,$
$\lambda_{4}(L_{0,i},L_{\alpha,k})=\lambda_{4}(L_{1,j},L_{\alpha,k})=0\,(k\neq 0,\,i,j\neq 0).$
When $k=0$ or $i,j=0$, we have $k(1-\delta_{i,0})=k(1-\delta_{j,0})=0$, thus $\Delta(L_{\alpha,k})=0$. ∎
Lemma 9.
Let $\Delta$ be a $2$-local derivation on $\mathcal{L}(\Gamma)$ such that $\Delta(L_{\alpha,i})=0$. Then for any $x=\sum_{\begin{subarray}{c}\gamma\in\Gamma\\ k\in\mathbb{Z}\end{subarray}}(m_{\gamma,k}L_{\gamma,k}+n_{\gamma,k}H_{\gamma,k})\in\mathcal{L}(\Gamma)$,
$\Delta(x)=0.$
Proof.
Given that $\Delta(L_{\alpha,i})=0$, by Lemma 8, we have
$\Delta(x)$ $=$ $\Delta_{L_{0,i},x}(x)$
$=$ $[a_{0,j}(L_{0,i},x)L_{0,j}+b_{0,j}(L_{0,i},x)H_{0,j},x]+\lambda_{3}(L_{0,i},x)D_{b}(x)$
$=$ $\sum_{\begin{subarray}{c}\gamma\in\Gamma k\in\mathbb{Z}\end{subarray}}(-\gamma m_{\gamma,k}a_{0,j}L_{\gamma,j+k}-\gamma n_{\gamma,k}a_{0,j}(L_{0,i},x)H_{\gamma,j+k})+n_{\gamma,k}b\lambda_{3}(L_{0,i},x)(H_{\gamma,k})$,
$\Delta(x)$ $=$ $\Delta_{L_{\alpha,i},x}(x)$
$=$ $[a_{\alpha,j}(L_{\alpha,i},x)L_{\alpha,j}+b_{0,j}(L_{\alpha,i},x)H_{0,j},x]+(1-\delta_{\alpha,0})\lambda_{1}(L_{\alpha,i},x)D_{f}(x)$
$+(1-\delta_{\alpha,0})\lambda_{2}(L_{\alpha,i},x)D_{g}(x)+\lambda_{3}(L_{\alpha,i},x)D_{b}(x)+(1-\delta_{i,0})\lambda_{4}(L_{\alpha,i},x)D^{\rho}(x)$
$=$ $\sum_{\begin{subarray}{c}\gamma\in\Gamma\\ k\in\mathbb{Z}\end{subarray}}((\alpha-\gamma)m_{\gamma,k}a_{\alpha,j}(L_{\alpha,i},x)L_{\alpha+\gamma,j+k}-\gamma n_{\gamma,k}a_{\alpha,j}(L_{\alpha,i},x)H_{\alpha+\gamma,j+k})$
$+f_{\gamma}(1-\delta_{\alpha,0})\lambda_{1}(L_{\alpha,i},x)(m_{\gamma,k}L_{\gamma,k}+n_{\gamma,k}H_{\gamma,k})+g_{\alpha}(1-\delta_{\alpha,0})\lambda_{2}(L_{\alpha,i},x)(m_{\gamma,k}L_{\gamma,k})$
$+b\lambda_{3}(L_{\alpha,i},x)(n_{\gamma,k}H_{\gamma,k})+k(1-\delta_{i,0})\lambda_{4}(L_{\alpha,i},x)(m_{\gamma,k}L_{\gamma,k-1}+n_{\gamma,k}H_{\gamma,k-1})$.
If there exists $m_{\alpha,i}\neq 0$, taking sufficiently many $\alpha\in\Gamma,\,i\in\mathbb{Z}$, we have $a_{\alpha,j}(L_{\alpha,i},x)=a_{0,j}(L_{\alpha,i},x)=0,\,i,j\in\mathbb{Z}$. Similarly $\lambda_{3}(L_{0,i},x)=\lambda_{3}(L_{\alpha,i},x)=0$, and $\lambda_{1}(L_{\alpha,i},x)=\lambda_{2}(L_{\alpha,i},x)=0(\alpha\neq 0),\,\lambda_{4}(L_{\alpha,i},x)=0(i\neq 0)$. When $\alpha=0$, we have $f_{\alpha}(1-\delta_{\alpha,0})$ and $g_{\alpha}(1-\delta_{\alpha,0})=0$. When $i=0$, we have $i(1-\delta_{i,0})=0$. ∎
Theorem 10.
Every $2$-local derivation on $\mathcal{L}(\Gamma)$ is a derivation.
Proof.
Let $\Delta$ be a $2$-local derivation on $\mathcal{L}(\Gamma)$. Let $\Delta_{L_{0,i},L_{1,j}}$ be a derivation on $\mathcal{L}(\Gamma)$ satisfies
$\Delta(L_{0,i})$ $=$ $\Delta_{L_{0,i},L_{1,j}}(L_{0,i})$,
$\Delta(L_{1,j})$ $=$ $\Delta_{L_{0,i},L_{1,j}}(L_{1,j})$.
Let $\Delta(1)=\Delta-\Delta_{L_{0,i},L_{1,j}}$. Then $\Delta(1)$ is a $2$-local derivation satisfying
$\Delta(1)(L_{0,i})=\Delta(1)(L_{1,j})=0.$
By Lemma 8, we have $\Delta(1)(L_{\alpha,k})=0$ and by Lemma 9, we have $\Delta(1)(x)=0,\,\mbox{for all}\,x\in\mathcal{L}(\Gamma)$. Then $\Delta(1)=\Delta-\Delta_{L_{0,i},L_{1,j}}\equiv 0$. Furthermore, $\Delta=\Delta_{L_{0,i},L_{1,j}}$ is a derivation on $\mathcal{L}(\Gamma)$. ∎
4. Biderivations on $\mathcal{L}(\Gamma)$
4.1. biderivations on $\mathcal{L}(\Gamma)$
Recall the definitions of biderivations and inner biderivations. Let $L$ be an algebra and $f$ be a bilinear map on $L\times L\rightarrow L$. If it is a derivation with respect to both components, meaning that for all $x\in L$, there exists a linear map from $L$ to itself $\phi_{x},\psi_{x}$, and $\phi_{x}=f(x,\cdot),\psi_{x}=f(\cdot,x)$ are derivations on $L$, satisfying that
$f(x\circ y,z)$ $=$ $x\circ f(y,z)+f(x,z)\circ y,$
$f(x,y\circ z)$ $=$ $f(x,y)\circ z+y\circ f(x,z).$
Then $f$ is called a biderivation of $L$. A biderivation $f$ that satisfies $f(x,y)=\lambda[x,y],$ where $\lambda\in\mathbb{F}$, is called an inner biderivation.
Lemma 11.
$\mathcal{L}(\Gamma)$ is perfect.
Proof.
Since [,][,][ , ] is bilinear, we only need to verify
$L_{\alpha,i}$ $=$ $\frac{1}{\alpha}[L_{\alpha-\beta,i-j},L_{\beta,j}]\quad(\alpha\neq 0,\alpha,\beta\in\Gamma,i,j\in\mathbb{Z}),$
$L_{0,i}$ $=$ $\frac{1}{2}[L_{1,i-j},L_{-1,j}]\quad(\alpha=0,i,j\in\mathbb{Z}),$
$H_{\alpha,i}$ $=$ $-\frac{1}{\beta}[L_{\alpha-\beta,i-j},H_{\beta,j}]\quad(\beta\neq 0,\alpha,\beta\in\Gamma,i,j\in\mathbb{Z}),$
$H_{0,i}$ $=$ $[L_{1,i-j},H_{-1,j}]\quad(\beta=0,i,j\in\mathbb{Z}).$
Then (Γ)=[$\mathcal{L}(\Gamma)=[\mathcal{L}(\Gamma),\mathcal{L}(\Gamma)]$. ∎
Lemma 12.
(see [11]) Let $L$ be perfect algebra and $f$ be a biderivation on $L$. If $\alpha\in C(L)$, then for all $x\in L$, $f(x,\alpha)=f(\alpha,x)=0$.
Lemma 13.
(see [12]) Let $f$ be a biderivation on $L$. Then
$[f(x,y),[u,v]]=[[x,y],f(u,v)],\,x,y,u,v\in L.$
In particular, $[f(x,y),[x,y]]=0,\,x,y\in L$.
Lemma 14.
(see [12]) Let $f$ be a biderivation on $L$. If $[x,y]=0,\,x,y\in L$, then $f(x,y)\in C(L)$.
Theorem 15.
Every biderivation of $\mathcal{L}(\Gamma)$ is inner.
Proof.
Let $f$ be biderivation on $\mathcal{L}(\Gamma)$.
(i) There exists $\lambda\in\mathbb{F}$ such that $f(L_{0,i},L_{m,j})\equiv\lambda[L_{0,i},L_{m,j}](\mathrm{mod}\mathbb{F}H_{0,i+j}),\,m\in\Gamma,i,j\in\mathbb{Z}.$
If $m=0$, then $[L_{0,i},L_{0,j}]=0$, by Lemma 15 we conclude that
$f(L_{0,i},L_{0,j})\in C(\mathcal{L}(\Gamma))=\mathbb{F}H_{0,i+j}.$
If $m\neq 0$, let
$f(L_{0,i},L_{m,j})=\sum_{\begin{subarray}{c}n\in\Gamma\\ i,j\in\mathbb{Z}\end{subarray}}(a_{n}L_{n,i+j}+b_{n}H_{n,i+j}).$
By Lemma 14, we conclude
$[[L_{0,i},L_{m,j}],f(L_{0,i},L_{m,j})]=0.$
Since
$[-mL_{m,i+j},a_{n}L_{n,i+j}+b_{n}H_{n,i+j}]=0,\,n\in\Gamma,i,j\in\mathbb{Z},$
then it follows that
$-ma_{n}(m-n)=0,\quad mnb_{n}=0.$
Furthermore, $a_{n}=0(m\neq n),\,b_{n}=0(n\neq 0)$, we conclude that
$f(L_{0,i},L_{m,j})$$\equiv$ $0(\mathrm{mod}\mathbb{F}H_{0,i+j}),\,m,n\neq 0,$
$f(L_{0,i},L_{m,j})$$\equiv$ $a_{m}L_{m,i+j}(\mathrm{mod}\mathbb{F}H_{0,i+j}),\,m\neq n,n=0.$
Thus $f(L_{0,i},L_{m,j})\equiv\lambda[L_{0,i},L_{m,j}](\mathrm{mod}\mathbb{F}H_{0,i+j}),\,m\in\Gamma,i,j\in\mathbb{Z}$.
(ii)$f(L_{0,i},H_{m,j})\equiv\lambda[L_{0,i},H_{m,j}](\mathrm{mod}\mathbb{F}H_{0,i+j}),\,m\in\Gamma,i,j\in\mathbb{Z}$.
If $m=0$, $H_{0,j}\in C(\mathcal{L}(\Gamma))$, then $[L_{0,i},H_{0,j}]=0$, it follows that $f(L_{0,i},H_{0,j})\in C(\mathcal{L}(\Gamma))$.
When $m\neq 0$, let
$f(L_{0,i},H_{m,j})=\sum_{\begin{subarray}{c}n\in\Gamma\\ i,j\in\mathbb{Z}\end{subarray}}(a_{n}L_{n,i+j}+b_{n}H_{n,i+j}).$
Since
$[[L_{0,j},H_{m,j}],f(L_{0,j},H_{m,j})]=0,$
then we have
$[-mH_{m,i+j},a_{n}L_{n,i+j}+b_{n}H_{n,i+j}]=0,\,n\in\Gamma,i,j\in\mathbb{Z}.$
We conclude $-m^{2}a_{n}=0$, thus $a_{n}=0$, then
$f(L_{0,i},H_{m,j})\equiv\sum_{\begin{subarray}{c}n\in\Gamma\\ i,j\in\mathbb{Z}\end{subarray}}b_{n}H_{n,i+j}.$
By
$[f(L_{0,i},H_{m,j}),[L_{0,i},L_{1,j}]]=[[L_{0,i},H_{m,j}],f(L_{0,i},L_{1,j})],$
we have
$[b_{n}H_{n,i+j},-L_{1,i+j}]=[-mH_{m,i+j},\lambda L_{1,i+j}],$
which says
$-nb_{n}H_{n+1,2(i+j)}=-m\lambda H_{m+1,2(i+j)}.$
We have $m=n$, which gives $f(L_{0,i},H_{m,j})\equiv\lambda[L_{0,i},H_{m,j}](\mathrm{mod}\mathbb{F}H_{0,i+j}),\,m\in\Gamma,i,j\in\mathbb{Z}$.
(iii)$f(L_{0,i},x)\equiv\lambda[L_{0,i},x](\mathrm{mod}\mathbb{F}H_{0,i+j}),\,x\in\mathcal{L}(\Gamma),i,j\in\mathbb{Z}$.
Let
$x=\sum_{\begin{subarray}{c}m\in\Gamma\\ i,j\in\mathbb{Z}\end{subarray}}(a_{m}L_{m,j}+b_{m}H_{m,j}).$
We have
$f(L_{0,i},x)$ $=$ $\sum_{\begin{subarray}{c}m\in\Gamma\\ i,j\in\mathbb{Z}\end{subarray}}\left\{a_{m}f(L_{0,i},L_{m,j})+b_{m}f(L_{0,i},H_{m,j})\right\}$
$=$ $\sum_{\begin{subarray}{c}m\in\Gamma\\ i,j\in\mathbb{Z}\end{subarray}}\left\{a_{m}\lambda[L_{0,i},L_{m,j}]+b_{m}\lambda[L_{0,i},H_{m,j}]\right\}(\mathrm{mod}\mathbb{F}H_{0,i+j})$
$=$ $\lambda[L_{0,i},\sum_{\begin{subarray}{c}m\in\Gamma\\ j\in\mathbb{Z}\end{subarray}}(a_{m}L_{m,j}+b_{m}H_{m,j})](\mathrm{mod}\mathbb{F}H_{0,i+j})$
$=$ $\lambda[L_{0,i},x](\mathrm{mod}\mathbb{F}H_{0,i+j}).$
which gives $f(L_{0,i},x)\equiv\lambda[L_{0,i},x](\mathrm{mod}\mathbb{F}H_{0,i+j}),x\in\mathcal{L}(\Gamma),i,j\in\mathbb{Z}$.
(iv)$f(x,y)\equiv\lambda[x,y](\mathrm{mod}\mathbb{F}H_{0,i+j}),\,x,y\in\mathcal{L}(\Gamma),i,j\in\mathbb{Z}$.
Following
$[f(x,y),[L_{0,i},z]]=[[x,y],f(L_{0,i},z)],\,z\in\mathcal{L}(\Gamma),$
we have
$[f(x,y)-\lambda[x,y],[L_{0,i},z]]=0,\,z\in\mathcal{L}(\Gamma),$
which gives $f(x,y)-\lambda[x,y]\in C(\mathcal{L}(\Gamma))$, thus $f(x,y)\equiv\lambda[x,y](\mathrm{mod}\mathbb{F}H_{0,i+j})$.
(v)$f(x,y)\equiv\lambda[x,y],\,x,y\in\mathcal{L}(\Gamma)$.
Assume that
$f(x,y)=\lambda[x,y]+\alpha(x,y)H_{0,i+j},$
where $\alpha$ is a bilinear map on $\mathcal{L}(\Gamma)\times\mathcal{L}(\Gamma)\rightarrow\mathbb{F}$. By Lemma 13
$f([x,y],z)$ $=$ $\lambda[[x,y],z]+\alpha([x,y],z)H_{0,i+j},$
$f([x,y],z)$ $=$ $[x,\lambda[y,z]+\alpha(y,z)H_{0,i+j}]+[\lambda[x,z]+\alpha(x,z)H_{0,i+j},y]$
$=$ $\lambda([x,[y,z]]+\lambda[[x,z],y]).$
We have
$\alpha([x,y],z)=0,\,x,y,z\in\mathcal{L}(\Gamma).$
Since $\mathcal{L}(\Gamma)$ is perfect, we conclude $\alpha(x,y)=0,\,x,y\in\mathcal{L}(\Gamma)$. Finally, we have that $f(x,y)\equiv\lambda[x,y],\,\forall x,y\in\mathcal{L}(\Gamma)$. That is, $f$ is an inner derivation on $\mathcal{L}(\Gamma)$. ∎
4.2. Applications
Recall the definition of linear commuting maps. If a linear map $\phi$ on Lie algebra $L$, which subjects to $[\phi(x),x]=0,\,\forall x\in L$, then $\phi$ called a linear commuting map on $L$. If $f(x,y)=[\phi(x),y]=[x,\phi(y)]$, then $f$ is a biderivation on $L$. (see [12])
Theorem 16.
Any linear map on $\mathcal{L}(\Gamma)$ is commuting if and only if there exists $\lambda\in\mathbb{F}$ and a linear map $\tau:\,\mathcal{L}(\Gamma)\rightarrow C(\mathcal{L}(\Gamma))$ such that $\phi(x)=\lambda x+\tau(x),\,x\in\mathcal{L}(\Gamma)$.
Proof.
Since
$[\phi(x),x]$ $=$ $[\lambda x+\tau(x),x]$
$=$ $\lambda[x,x]+[\tau(x),x]$
$=$ $0$
Then $\phi$ is a linear commuting map.
Conversely, let $f(x,y)=[\phi(x),y],\,x,y\in\mathcal{L}(\Gamma)$. Then $f$ is a biderivation on $\mathcal{L}(\Gamma)$. By Theorem 15, we have
$f(x,y)=[\phi(x),y]=\lambda[x,y].$
Since $\phi(x)-\lambda x\in C(\mathcal{L}(\Gamma))$, then we have a linear map $\tau:\mathcal{L}(\Gamma)\rightarrow C(\mathcal{L}(\Gamma))$ such that $\phi(x)=\lambda x+\tau(x),\,x\in\mathcal{L}(\Gamma)$. ∎
[…]
$x\cdot y=y\cdot x,$
$[x,y]\cdot z=x\cdot(y\cdot z)-y\cdot(x\cdot z),$
$x\cdot[y,z]=[x\cdot y,z]+[y,x\cdot z],\,x,y,z\in L.$
We also call $(\mathcal{L}(\Gamma),[,],\cdot)$ is a commutative $post$-Lie algebra (see [13]).
Lemma 17.
(see, [13]) Let $(L,[,],\cdot)$ be a commutative $post$-Lie algebra. If $f:L\times L\rightarrow L$ subjects to $f(x,y)=x\cdot y,\,x,y\in L$, then $f$ is a biderivation on $L$.
Theorem 18.
Any commutative $post$-Lie algebra structure on $\mathcal{L}(\Gamma)$ is trivial.
Proof.
Assume $(\mathcal{L}(\Gamma),[,],\cdot)$ is a commutative $post$-Lie algebra. By Theorem 15 and Lemma 17, we know there exists $\lambda\in\mathbb{C}$ such that
$f(x,y)=x\cdot y=\lambda[x,y].$
Since $x\cdot y=y\cdot x$, we have
$\lambda[x,y]=\lambda[y,x].$
Thus $02\lambda[x,y]=$ 0, we conclude $\lambda=0$. That is $f(x,y)=x\cdot y=0$. ∎
5. Automorphism groups on $\mathcal{L}(\Gamma)$
Let $L=\mathrm{span}_{\mathbb{F}}\{L_{\alpha,i}\,|\,\alpha\in\Gamma,i\in\mathbb{Z}\},\,H=\mathrm{span}_{\mathbb{F}}\{H_{\alpha,i}\,|\,\alpha\in\Gamma,i\in\mathbb{Z}\}.$
Lemma 19.
$I=\left\{H_{\alpha,i}\,|\,\alpha\in\Gamma,i\in\mathbb{Z}\right\}$ is a unique maximal ideal on $\mathcal{L}(\Gamma)$.
Proof.
By the definition of $\mathcal{L}(\Gamma)$, $I$ is an ideal of $\mathcal{L}(\Gamma)$. Since $\mathcal{L}(\Gamma)/I$ is isomorphic to the generalized centerless Loop Virasoro algebra. The centerless Loop Virasoro algebra is a simple Lie algebra, then $I$ is a maximal ideal. Next, we prove $I=I^{\prime}$. Let $I^{\prime}$ be another maximal ideal of $\mathcal{L}(\Gamma)$. For all $0\neq x\in I^{\prime}$, let $x=l+h$, where $l\in L,\,h\in H$. If $l\neq 0$, then $I^{\prime}=\mathcal{L}(\Gamma)$, which is a contradiction. Thus $l=0$. ∎
Lemma 20.
Let $\mathcal{LV}(\Gamma)$ and $\mathcal{LV}(\Gamma^{\prime})$ be two generalized centerless Loop Virasoro algebra. $\mathcal{LV}(\Gamma)\cong\mathcal{LV}(\Gamma^{\prime})$ if and only if there there exists $a\in\mathbb{F}^{\ast}$, such that $a\Gamma^{\prime}=\Gamma$, where $\mathbb{F}^{\ast}=\mathbb{F}\setminus\{0\}$.
Proof.
Notice that the ”only if” part is obvious. We now prove the ”if” part.
Given that $\mathcal{LV}(\Gamma)\cong\mathcal{LV}(\Gamma^{\prime})$, let $\theta:\mathcal{LV}(\Gamma)\rightarrow\mathcal{LV}(\Gamma^{\prime})$ be a Lie algebra isomorphism. Since $\mathbb{F}L_{0,0},\,\mathbb{F}L^{\prime}_{0,0}$ are the Cartan subalgebras of $\mathcal{LV}(\Gamma)$ and $\mathcal{LV}(\Gamma^{\prime})$, respectively, then for all $a\in\mathbb{F}^{\ast}$ such that $\theta(L_{0,0})=aL^{\prime}_{0,0}.$
$\theta[L_{0,0},L_{\alpha,j}]$ $=$ $\theta(-\alpha L_{\alpha,j})$
$=$ $-\alpha\theta(L_{\alpha,j}),$
$\theta[L_{0,0},L_{\alpha,j}]$ $=$ $[\theta(L_{0,0}),\theta(L_{\alpha,j})]$
$=$ $a[L^{\prime}_{0,0},\theta(L_{\alpha,j})].$
Thus $[L^{\prime}_{0,0},\theta(L_{\alpha,j})]=-a^{-1}\alpha\theta(L_{\alpha,j})$, that is $-a^{-1}\alpha\in\Gamma^{\prime}$, we conclude $\Gamma\subset a\Gamma^{\prime}$. Similarly, it can be proved that $a\Gamma^{\prime}\subset\Gamma$. ∎
Lemma 21.
$\mathcal{L}(\Gamma)\cong\mathcal{L}(\Gamma^{\prime})$ if and only if there exists $a\in\mathbb{F}^{\ast}$, such that $a\Gamma^{\prime}=\Gamma$.
Proof.
Let $\theta:\mathcal{L}(\Gamma)\rightarrow\mathcal{L}(\Gamma^{\prime})$ be a Lie algebra isomorphism and $I,I^{\prime}$ be the maximal ideals of $\mathcal{L}(\Gamma)$ and $\mathcal{L}(\Gamma^{\prime})$, respectively. By Lemma 19, we have $\theta(I)=I^{\prime}$, and by Lemma 20, we have $\mathcal{L}(\Gamma)/I\cong\mathcal{L}(\Gamma^{\prime})/I^{\prime}$, then exists $a\in\mathbb{F}^{\ast}$.
Conversely, if $a\Gamma^{\prime}=\Gamma$, define a mapping by
$\theta:\mathcal{L}(\Gamma)$$\rightarrow$ $\mathcal{L}(\Gamma^{\prime})$
[…]
Then $\theta$ is an isomorphism obviously. ∎
For all $\theta\in\mathrm{Aut}\mathcal{L}(\Gamma)$, considering $\theta$ act on the basis of $\mathcal{L}(\Gamma)$, for all $\alpha\in\Gamma,\,i\in\mathbb{Z}$, assume $f_{\alpha,i}\in\mathbb{F}[t^{1},t^{-1}]\in H,$
$\theta(L_{\alpha,i})$ $=$ $L_{\frac{\alpha}{a}}f_{\alpha,i},$ (5.1)
$\theta(H_{\alpha,i})$ $=$ $H_{\frac{\alpha}{a}}f_{\alpha,i}.$ (5.2)
Thus
$\theta([L_{\alpha,i},L_{\beta,j}])$ $=$ $\theta((\alpha-\beta)L_{\alpha+\beta,i+j})$
$=$ $(\alpha-\beta)L_{\frac{\alpha+\beta}{a}}f_{\alpha+\beta,i+j},$
$\theta([L_{\alpha,i},L_{\beta,j}])$ $=$ $[\theta(L_{\alpha,i}),\theta(L_{\beta,j})]$
$=$ $[L_{\frac{\alpha}{a}}f_{\alpha,i},L_{\frac{\beta}{a}}f_{\beta,j}]$
$=$ $\frac{\alpha-\beta}{a}L_{\frac{\alpha+\beta}{a}}f_{\alpha,i}f_{\beta,j},$
$\theta([L_{\alpha,i},H_{\beta,j}])$ $=$ $\theta(-\beta H_{\alpha+\beta,i+j})$
$=$ $-\beta H_{\frac{\alpha+\beta}{a}}f_{\alpha+\beta,i+j},$
$\theta([L_{\alpha,i},H_{\beta,j}])$ $=$ $[\theta(L_{\alpha,i}),\theta(H_{\beta,j})]$
$=$ $[L_{\frac{\alpha}{a}}f_{\alpha,i},H_{\frac{\beta}{a}}f_{\beta,j}]$
$=$ $-\frac{\beta}{a}H_{\frac{\alpha+\beta}{a}}f_{\alpha,i}f_{\beta,j}.$
By comparing the expressions, we have:
$a(\alpha-\beta)f_{\alpha+\beta,i+j}=(\alpha-\beta)f_{\alpha,i}f_{\beta,j},$
$a\beta f_{\alpha+\beta,i+j}=\beta f_{\alpha,i}f_{\beta,j},\,\forall\alpha,\beta\in\Gamma,\,i,j\in\mathbb{Z}.$
Lemma 22.
Let $a\in\mathbb{F}^{\ast},\,f_{\alpha},f_{\beta}\in\mathbb{F}[t^{1},t^{-1}]$. For all $\alpha,\beta\in\Gamma,\,i,j\in\mathbb{Z}$, then $af_{\alpha+\beta,i+j}=f_{\alpha,i}f_{\beta,j}$.
Proof.
The proof is similar to that in Lemma 3 and Lemma 4. ∎
Let $\mathbb{F}[t^{1},t^{-1}]^{\ast}=\{at^{i}\,|\,a\in\mathbb{F}^{\ast},i\in\mathbb{Z}\}$. By Lemma 22, we have $f_{0,0}$ is a non-zero element on $\mathbb{F}[t^{1},t^{-1}]$, then $f_{0,0}\in\mathbb{F}[t^{1},t^{-1}]^{\ast}$. Thus, we can write $f_{\alpha,i}$ in the following form:
$f_{\alpha,i}:=a\mu(\alpha,i)t^{\epsilon(\alpha,i)},\,\forall\alpha\in\Gamma,\,i\in\mathbb{Z},$
where $\mu(\alpha,i)\in\mathbb{F}^{\ast},\,\epsilon(\alpha,i)\in\mathbb{Z}$, thus for all $\alpha_{1},\alpha_{2}\in\Gamma,\,i,j\in\mathbb{Z}$,
$\mu(\alpha_{1},i)\mu(\alpha_{2},j)=\mu(\alpha_{1}+\alpha_{2},i+j),$
$\epsilon(\alpha_{1},i)+\epsilon(\alpha_{2},j)=\epsilon(\alpha_{1}+\alpha_{2},i+j).$
Then, (5.1) and (5.2) can be written as follows:
$\theta(L_{\alpha,i})$ $=$ $a\mu(\alpha,i)L_{{\frac{\alpha}{a}},\epsilon(\alpha,i)},$ (5.3)
$\theta(H_{\alpha,i})$ $=$ $a\mu(\alpha,i)H_{{\frac{\alpha}{a}},\epsilon(\alpha,i)}.$ (5.4)
Based on the above formulas, we obtain the following automorphisms on $\mathcal{L}(\Gamma)$.
Lemma 23.
(i) For all $a\in A=\{a\in\mathbb{F}^{\ast}\,|\,a\Gamma=\Gamma\}$, the map
$\theta_{a}:\mathcal{L}(\Gamma)$$\longrightarrow$ $\mathcal{L}(\Gamma)$
[…]
is an automorphism on $\mathcal{L}(\Gamma)$.
(ii) For all $\phi\in\mathrm{Hom}(\Gamma,\mathbb{Z})$, the map
$\theta_{\phi}:\mathcal{L}(\Gamma)$$\longrightarrow$ $\mathcal{L}(\Gamma)$
[…]
is an automorphism on $\mathcal{L}(\Gamma)$.
(iii) For all $\chi\in\chi(\Gamma)$, $\chi(\Gamma)$ is the set consisting of group homomorphisms from $\Gamma$ to $\mathbb{F}^{\ast}$, the map
$\theta_{\chi}:\mathcal{L}(\Gamma)$$\longrightarrow$ $\mathcal{L}(\Gamma)$
[…]
is an automorphism on $\mathcal{L}(\Gamma)$.
(iv) For all $\psi\in\mathrm{Aut}\mathbb{Z}=\{id,-id\}$, the map
$\theta_{\psi}:\mathcal{L}(\Gamma)$$\longrightarrow$ $\mathcal{L}(\Gamma)$
[…]
is an automorphism on $\mathcal{L}(\Gamma)$.
(v) Given $b\in\mathbb{F}^{\ast}$, the map
$\theta_{b}:\mathcal{L}(\Gamma)$$\longrightarrow$ $\mathcal{L}(\Gamma)$
[…]
is an automorphism on $\mathcal{L}(\Gamma)$.
Proof.
Obviously. ∎
Recall the [7], we have the set $(A\times\mathrm{Hom}(\Gamma,\mathbb{Z})\times\chi(\Gamma)\times\mathrm{Aut}\mathbb{Z}\times\mathbb{F^{\ast}})$ is a multiplicative group for which $(1,0,0,1,1)$ is the identify element, it satisfies
$(a_{1},\phi_{1},\chi_{1},\psi_{1},b_{1})\cdot(a_{2},\phi_{2},\chi_{2},\psi_{2},b_{2})=(a_{1}a_{2},\phi_{1}\iota_{a^{-1}_{2}}+\psi_{1}\phi_{2},b^{\phi_{2}}_{1}+\chi_{2}+\chi_{1}\iota_{a_{2}^{-1}},\psi_{1}\psi_{2},b^{\psi_{2}}_{1}b_{2}),$
where $\iota_{a}$ is an automorphism on $\Gamma$, satisfying $\iota_{a}(\alpha)=a\alpha$. If $b^{\phi}\in\chi(\Gamma)$, it satisfies $(b^{\phi})(\alpha)=b^{\phi(\alpha)}$, where $\alpha\in\Gamma,\,b\in\mathbb{F}^{\ast},\,\phi\in\mathrm{Hom}(\Gamma,\mathbb{Z}).$
Theorem 24.
$\mathrm{Aut}\mathcal{L}(\Gamma)\cong(A\times Hom(\Gamma,\mathbb{Z})\times\chi(\Gamma)\times\mathrm{Aut}\mathbb{Z}\times\mathbb{F}^{\ast})$.
Proof.
Let $\Psi$ be a map from $A\times\mathrm{Hom}(\Gamma,\mathbb{Z})\times\chi(\Gamma)\times\mathrm{Aut}\mathbb{Z}\times\mathbb{F}^{\ast}$ to $\mathrm{Aut}\mathcal{L}(\Gamma)$,
$\Psi:A\times\mathrm{Hom}(\Gamma,\mathbb{Z})\times\chi(\Gamma)\times\mathrm{Aut}\mathbb{Z}\times\mathbb{F}^{\ast}$$\longrightarrow$ $\mathrm{Aut}\mathcal{L}(\Gamma)$
[…]
Now we prove $\Psi$ is an isomorphism. For all
$(a_{1},\phi_{1},\chi_{1},\psi_{1},b_{1}),(a_{2},\phi_{2},\chi_{2},\psi_{2},b_{2})\in A\times\mathrm{Hom}(\Gamma,\mathbb{Z})\times\chi(\Gamma)\times\mathrm{Aut}\mathbb{Z}\times\mathbb{F}^{\ast},$
on the one hand,
$\Psi(a_{1},\phi_{1},\chi_{1},\psi_{1},b_{1})\Psi(a_{2},\phi_{2},\chi_{2},\psi_{2},b_{2})(L_{\alpha,i})$
$=$ $\Psi(a_{1},\phi_{1},\chi_{1},\psi_{1},b_{1})(\theta_{a_{2}}\theta_{\phi_{2}}\theta_{\chi_{2}}\theta_{\psi_{2}}\theta_{b_{2}})(L_{\alpha,i})$
$=$ $\Psi(a_{1},\phi_{1},\chi_{1},\psi_{1},b_{1})(\theta_{a_{2}}\theta_{\phi_{2}}\theta_{\chi_{2}}\theta_{\psi_{2}})(b_{2}^{i}L_{\alpha,i})$
$=$ $\Psi(a_{1},\phi_{1},\chi_{1},\psi_{1},b_{1})(\theta_{a_{2}}\theta_{\phi_{2}}\theta_{\chi_{2}})(b^{i}_{2}L_{\alpha,\psi_{2}(i)})$
$=$ $\Psi(a_{1},\phi_{1},\chi_{1},\psi_{1},b_{1})(\theta_{a_{2}}\theta_{\phi_{2}})(b_{2}^{i}\chi_{2}(\alpha)L_{\alpha,\psi_{2}(i)})$
$=$ $\Psi(a_{1},\phi_{1},\chi_{1},\psi_{1},b_{1})(\theta_{a_{2}})(b_{2}^{i}\chi_{2}(\alpha)L_{\alpha,\psi_{2}(i)+\phi_{2}(\alpha)})$
$=$ $\Psi(a_{1},\phi_{1},\chi_{1},\psi_{1},b_{1})(a_{2}b_{2}^{i}\chi_{2}(\alpha)L_{\frac{\alpha}{a_{2}},\psi_{2}(i)+\phi_{2}(\alpha)})$
$=$ $(\theta_{a_{1}}\theta_{\phi_{1}}\theta_{\chi_{1}}\theta_{\psi_{1}}\theta_{b_{1}})(a_{2}b_{2}^{i}\chi_{2}(\alpha)L_{\frac{\alpha}{a_{2}},\psi_{2}(i)+\phi_{2}(\alpha)})$
$=$ $(\theta_{a_{1}}\theta_{\phi_{1}}\theta_{\chi_{1}}\theta_{\psi_{1}})(a_{2}b_{1}^{\psi_{2}(i)+\phi_{2}(\alpha)}b_{2}^{i}\chi_{2}(\alpha)L_{\frac{\alpha}{a_{2}},\psi_{2}(i)+\phi_{2}(\alpha)})$
$=$ $(\theta_{a_{1}}\theta_{\phi_{1}}\theta_{\chi_{1}})(a_{2}b_{1}^{\psi_{2}(i)+\phi_{2}(\alpha)}b_{2}^{i}\chi_{2}(\alpha)L_{\frac{\alpha}{a_{2}},\psi_{1}(\psi_{2}(i)+\phi_{2}(\alpha))})$
$=$ $(\theta_{a_{1}}\theta_{\phi_{1}})(a_{2}b_{1}^{\psi_{2}(i)+\phi_{2}(\alpha)}b_{2}^{i}\chi_{1}(\frac{\alpha}{a_{2}})\chi_{2}(\alpha)L_{\frac{\alpha}{a_{2}},\psi_{1}(\psi_{2}(i)+\phi_{2}(\alpha))})$
$=$ $(\theta_{a_{1}})(a_{2}b_{1}^{\psi_{2}(i)+\phi_{2}(\alpha)}b_{2}^{i}\chi_{1}(\frac{\alpha}{a_{2}})\chi_{2}(\alpha)L_{\frac{\alpha}{a_{2}},\psi_{1}(\psi_{2}(i)+\phi_{2}(\alpha))+\phi_{1}(\frac{\alpha}{a_{2}})})$
$=$ $a_{1}a_{2}b_{1}^{\psi_{2}(i)+\phi_{2}(\alpha)}b_{2}^{i}\chi_{1}(\frac{\alpha}{a_{2}})\chi_{2}(\alpha)L_{\frac{\alpha}{a_{1}a_{2}},\psi_{1}(\psi_{2}(i)+\phi_{2}(\alpha))+\phi_{1}(\frac{\alpha}{a_{2}})}.$
On the other hand,
$\Psi((a_{1},\phi_{1},\chi_{1},\psi_{1},b_{1})\cdot(a_{2},\phi_{2},\chi_{2},\psi_{2},b_{2}))(L_{\alpha,i})$
$=$ $\Psi(a_{1}a_{2},\phi_{1}\iota_{a^{-1}_{2}}+\psi_{1}\phi_{2},b^{\phi_{2}}_{1}+\chi_{2}+\chi_{1}\iota_{a_{2}^{-1}},\psi_{1}\psi_{2},b^{\psi_{2}}_{1}b_{2})(L_{\alpha,i})$
$=$ $a_{1}a_{2}(b_{1}^{\psi_{2}}b_{2})^{i}b_{1}^{\phi_{2}(\alpha)}\chi_{2}(\alpha)\chi(\frac{\alpha}{a_{2}})L_{\frac{\alpha}{a_{1}a_{2}},(\psi_{1}\psi_{2})(i)+(\phi_{1}\iota_{a^{-1}_{2}}+\psi_{1}\phi_{2})(\alpha)})$
$=$ $a_{1}a_{2}b_{1}^{\psi_{2}(i)+\phi_{2}(\alpha)}b_{2}^{i}\chi_{1}(\frac{\alpha}{a_{2}})\chi_{2}(\alpha)L_{\frac{\alpha}{a_{1}a_{2}},\psi_{1}(\psi_{2}(i)+\phi_{2}(\alpha))+\phi_{1}(\frac{\alpha}{a_{2}})}.$
Similarly, act on $H_{\alpha,i}$, we conclude that
$\Psi(a_{1},\phi_{1},\chi_{1},\psi_{1},b_{1})\Psi(a_{2},\phi_{2},\chi_{2},\psi_{2},b_{2})(H_{\alpha,i})=\Psi((a_{1},\phi_{1},\chi_{1},\psi_{1},b_{1})\cdot(a_{2},\phi_{2},\chi_{2},\psi_{2},b_{2}))(H_{\alpha,i}),$
thus $\Psi$ is a homomorphism. Next we prove $\Psi$ is a bijective map.
Let $\theta_{a_{1}}\theta_{\phi_{1}}\theta_{\chi_{1}}\theta_{\psi_{1}}\theta_{b_{1}}=\theta_{a_{2}}\theta_{\phi_{2}}\theta_{\chi_{2}}\theta_{\psi_{2}}\theta_{b_{2}}$. By $\theta_{a},\theta_{\phi},\theta_{\chi},\theta_{\psi},\theta_{b}$ are automorphisms on $\mathcal{L}(\Gamma)$, thus $\theta_{a_{1}}=\theta_{a_{2}},\theta_{\phi_{1}}=\theta_{\phi_{2}},\,\theta_{\chi_{1}}=\theta_{\chi_{2}},\,\theta_{\psi_{1}}=\theta_{\psi_{2}},\theta_{b_{1}}=\theta_{b_{2}}$, we conclude $(a_{1},\phi_{1},\chi_{1},\psi_{1},b_{1})=(a_{2},\phi_{2},\chi_{2},\psi_{2},b_{2})$, $\Psi$ is an injective map.
For all $\theta\in\mathrm{Aut}\mathcal{L}(\Gamma)$, by Lemma 21 and Lemma 22 we have $\theta(L_{\alpha,i})=a\mu(\alpha,i)L_{{\frac{\alpha}{a}},\epsilon(\alpha,i)},\theta(H_{\alpha,i})=a\mu(\alpha,i)H_{{\frac{\alpha}{a}},\epsilon(\alpha,i)}$. Let $\mu(\alpha,0)=\chi(\alpha),\,\mu(0,i)=b(i),\,\epsilon(\alpha,0)=\phi(\alpha),\,\epsilon(0,i)=\psi(i)$. Then $\Psi(a,\phi,\chi,\psi,b)=\theta$, therefore $\Psi$ is a surjective map. In conclusion, $\Psi$ is an isomorphism. ∎